OpenStudy (anonymous):

Simplify square root of 3 multiplied by the fifth root of 3.

OpenStudy (anonymous):

sqrt(3)*fifth_root(3) =3^(1/2) * 3^(1/5) =3^(1/2+1/5) =3^(7/10)

OpenStudy (anonymous):

thanks! <3 do you have time for another?

OpenStudy (anonymous):

sure :)

OpenStudy (anonymous):

Which of the following is equal to square root of the cube root of 6?

OpenStudy (anonymous):

6 1/6 6 1/3 6 2/3 6 3/2

OpenStudy (anonymous):

sqrt(cuberoot(6)) = (6^(1/3))^(1/2) = 6^(1/3 * 1/2) = 6^(1/6)

OpenStudy (anonymous):

6 1/6 ? :]

OpenStudy (anonymous):

yup :)

OpenStudy (anonymous):

thanks for helping :] if you have the time i have a couple more i need help with c:

OpenStudy (anonymous):

unfortunately i have to go soon, hopefully someone else can help you :(

OpenStudy (anonymous):

ah okay ..What is the exact value of the expression the square root of 294. - the square root of 24. + the square root 54.? Simplify if possible. 8the square root of 2. 8the square root of 6. 12the square root of 2. 12the square root of 6.

OpenStudy (anonymous):

@eashy just my last question :P

OpenStudy (anonymous):

sqrt(294)-sqrt(24)+sqrt(54) =sqrt(7*7*6)-sqrt(2*2*2*3)+sqrt(3*3*3*2) =7sqrt(6)-2sqrt(6)+3sqrt(6) =8sqrt(6)